OpenStudy (anonymous):

ok this is gonna seem much more simple than what u guys r doing but I havent done math in years, can u help? -4/13-1/2

OpenStudy (anonymous):

-4/13 = -18/26 -1/2 = -13/26 Go from there! Good luck!